OpenStudy (anonymous):

Phosphorus-32 has a half life of 14.3 days. how many days will it take for radioactive sample to decay to 1/8 it's size?

OpenStudy (anonymous):

iodine-131 has a half-life of 8 days. how many grams of an original 160 mg sample will remain after 40 days?

OpenStudy (jfraser):

How many half-lives have passes if you have 1/8 of what you started with? If you cut a piece of paper in half, you have 1/2 the paper left after 1 half-life. Cut that half in half again, and you have half of the half, or 1/4 remaining, after 2 half-lives. Cut that 1/4 in half again, and you have 1/2 of 1/4, or 1/8 of the original paper left, after 3 half-lives. In general, \[(\frac{1}{2})^n = (\frac{1}{8})\] so n must = 3. Three half-lives have passed. If each half-life takes 14.3 days, how long is 3 half-lives? The second question is the opposite of the first. If one half-live is 8 days, how many half-lives have passed after 40 days? Once you find that, cut the 160mg sample in half that many times.